Operations Report – December 17th

Crew 185
Operations Report Sunday 12/17/2017
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
The heater was not working properly, even after reseting it. Shannon adviced us to change the air filter.
– Generator :
The generator had a failure at 8:00pm, Fault message reading “Low battery”. Jump-start from Shannon’s Shuttle battery did not solve the issue. Generator had a second failure at 9:00pm, so we definitely turned it off for the night.
Generator (hours run):
Generator turned off at 9:00am.
Generator turned on at 5:00pm.
Generator had a failure at 8:00pm.
Jump-start done at 8:00pm.
Generator turned off at 9:00pm after a second failure.
Solar— SOC @ 5:00pm : 95%
Diesel – 80%
Propane – 62 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)
Water (trailer) – 300 gallons
Water (static) – 230 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATV’s Used: Yes
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:40 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  Yes
Hours: 00:40 Hours
Beginning charge: 100%
Ending charge: 90%
Currently charging: Yes
Opportunity rover used:  Yes
Hours: 00:40 Hours
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes
Curiosity rover used:  No
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: All Nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: Greenhab report is nominal and cleaned!
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

GreenHab Report – December 17th

Greenhab Report
David Murray
December 17th, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient with window/door open from 10:30 – ? (Someone closed door without notifying me)
Working Hour: 18:30
Outside temp during working hours: 1 C
Inside temperature high: 31 C
Inside temperature low: 15 C
Inside humidity: 16%
Functioning nominally
Turns on at 16 C and off at 20 C
Not functional
Shade Cloth On/Off
Average temperature:
Changes to Crops
Bean plants are no longer drooping
One bean plant has 9 flowers open or partially open
Two bean plants have one flower open or partially open
One bean plant has two flowers open or partially open
Dill weed has sprouted
Morning Research Observations
Daily Water Usage for Crops
5 gallons
Time(s) of Watering Crops
Around 13:00 and 18:30
Took inventory of all identifiable living plants in the greenhab.
Cucumber- 23
Tomato- 101
Beans- 31
Melon- 21
Pepper- 25
Many tomatoes need to be transplanted into larger pots
Support/Supplies Needed
Copyright © The Mars Society. All rights reserved. | Main Site